# A Survey of Vibe Coding with Large Language Models

### v1

@@ 6.2.1 AI-Native Development Interfaces
The integration of coding agents into the developer’s workspace has transformed the traditional notion of an Integrated Development Environment (IDE). Instead of passively editing code, developers now engage in a continuous dialogue with the system. AI-native development interfaces embed large language models directly within the editor, combining code completion, conversational reasoning, and contextual memory [84, 45, 1, 46].
Two dominant paradigms have emerged. The first emphasizes inline suggestion, where the system predicts the next tokens or statements directly within the editor, offering context-sensitive autocompletion and refactoring hints. The second paradigm emphasizes conversational interaction, in which the developer communicates with the agent using natural language to specify intent, request explanations, or perform debugging [468, 469, 470, 140, 471, 472].
Studies show that integrating both paradigms within a unified interface produces higher task efficiency and user satisfaction. Inline completion supports rapid iteration, while conversational interaction enhances explainability and collaboration [469, 471]. Emerging IDEs such as Cursor and Q Developer integrate contextual memory and multi-turn reasoning, maintaining persistent dialogue histories to provide more coherent assistance [141, 471, 473, 144].
@@ 6.2.2 Remote Development
The growth of cloud-native development practices has led to widespread adoption of remote and containerized development environments. Systems such as GitHub Codespaces [474] and DevContainer provide pre-configured, reproducible workspaces accessible via browsers or local editors. This paradigm allows coding agents to operate directly within standardized, resource-isolated environments synchronized with version control systems.
Remote development also supports computation-intensive workflows by allowing resource-demanding tasks to run on remote servers while the developer interacts through lightweight clients. SSH-based connections and browser-based terminals enable seamless transitions between local and cloud execution contexts. These capabilities are essential for agent-assisted programming, as they allow distributed collaboration while maintaining strict control over dependencies and runtime configurations [475].
@@ 6.2.3 Tool Integration Protocol Standards
Standardization efforts have become increasingly important for achieving interoperability between coding agents and existing developer tools. The Model Context Protocol (MCP) [96, 476] defines a universal interface for exchanging contextual information such as source code, documentation, and environment state. The Language Server Protocol (LSP) [142, 145] provides language-agnostic features including diagnostics, completion, and code navigation, while the Debug Adapter Protocol (DAP) [477, 478] standardizes debugging interactions.
Integrating these protocols allows coding agents to function as first-class participants within conventional development ecosystems. Agents can request, modify, and analyze project context seamlessly, while maintaining compatibility with existing IDEs and version control systems [479, 146, 143]. This protocol convergence promotes an open, extensible environment for hybrid human–AI programming collaboration [480, 481].
@@ 6.3.1 CI/CD Pipeline Integration
Continuous Integration and Continuous Deployment (CI/CD) frameworks are essential for maintaining the stability and quality of software systems. Within the context of coding agents, CI/CD processes ensure that automatically generated code passes through rigorous validation and testing stages before integration [155, 273, 482, 156]. Pipeline-as-Code practices enable reproducibility by representing build and deployment workflows as versioned artifacts [155, 483]. In cloud-native ecosystems, Kubernetes and Jenkins-based automation frameworks such as KubeSphere facilitate modular CI/CD pipelines. These pipelines automatically handle code building, testing, and deployment, while ensuring that agent-generated contributions are subject to the same review and validation processes as human-written code.
In the emerging paradigm of Vibe Coding, CI/CD pipelines are further extended with agent participation. Agents generate code, perform static analysis, and conduct self-testing, but human oversight remains indispensable for critical verification stages. Designing CI/CD architectures that integrate agents safely and transparently is a key challenge for reliable large-scale deployment [155].
@@ 6.3.2 Cloud Compute Orchestration
Distributed orchestration platforms employ frameworks dynamically provisioning computational resources with production deployments demonstrating scale operating on 25,000 CPU core clusters organized into specialized pod configurations [460]. The orchestration landscape has been enhanced by TOSCA (Topology and Orchestration Specification for Cloud Applications) providing interoperable models for describing cloud applications as typed directed topology graphs representing components as nodes and dependencies as links enabling developers to model cloud service structures with automated lifecycle management [157, 147, 148]. TOSCA’s platform-agnostic nature allows users to switch between cloud providers enabling applications to be migrated between vendors with minimal modification demonstrating strong compatibility with CI/CD technologies facilitating easier testing, re-deployment, and re-engineering [148, 149], with modern platforms like FogArm providing autonomous adaptation capabilities responding to application specification changes from CI/CD pipelines and infrastructural variations detected through distributed monitoring [157]. LLM-powered component integration adds complexity and opportunity where frameworks built on large language models, retrieval-augmented generation, and deep learning enable seamless integration with CI/CD pipelines supporting multi-format output generation and incorporating continuous learning for optimization [158]. Modern implementations configure CI/CD pipelines to build, test, and deploy applications within containerized environments enabling faster build times and more frequent releases with Continuous Deployment integrating directly with LLMOps to automatically deploy model updates based on performance monitoring enabling automated model transitioning, deployment, and monitoring while maintaining high reliability using GitOps pipelines ensuring automatic redeployment of updated agent code or model weights with Canary releases and A/B testing [159, 160].
@@ 6.3.3 Multi-Agent Collaboration Frameworks
Frameworks such as AutoGen, CrewAI, MetaGPT, and LangGraph exemplify this paradigm. AutoGen organizes heterogeneous agents that perform role-specific tasks such as requirement analysis, implementation, and unit testing [150, 151, 152]. CrewAI conceptualizes agent collaboration as a simulation of human teams, incorporating negotiation, task assignment, and feedback integration [153]. MetaGPT introduces a meta-planning layer, where high-level agents design task hierarchies and delegate execution to subordinate agents, thereby achieving structured decomposition [252, 154]. LangGraph formalizes collaboration as a graph of nodes and edges representing agents and communication pathways, which enables systematic reasoning over dependencies and task flows [484].
These frameworks collectively redefine how coding agents can scale across complex development pipelines. They provide fault tolerance by redistributing tasks when individual agents fail, improve modularity through role specialization, and support iterative verification through feedback loops among agents. Recent studies suggest that multi-agent collaboration enhances both the reliability and interpretability of large-scale code synthesis [485, 123].
Despite these advances, challenges remain regarding consistency, communication efficiency, and conflict resolution. Overly frequent message exchanges may lead to coordination overhead, while insufficient synchronization can cause semantic drift among agents. Ongoing research is exploring graph-structured communication protocols and shared reasoning buffers to improve coherence across collaborative agents [161, 486].
The development of multi-agent frameworks represents an essential direction for the evolution of Vibe Coding systems. These frameworks not only enhance scalability and adaptability but also embody a shift toward distributed intelligence in software engineering, where human developers and autonomous agents collaborate as peers within integrated ecosystems [487].
@@ 7.1.1 Syntax and Type Error Feedback
LLM-based coding agents rely on compiler feedback to iteratively improve generated code through error detection and repair cycles [488, 489, 1, 105, 117, 490]. Compiler feedback encompasses multiple granularities: coarse-grained feedback provides binary compilation success, while fine-grained feedback delivers detailed error reasons and locations [491]. Beyond basic compilation feedback, systems integrate test feedback from automated frameworks and static analysis from tools like pylint [492].
Modern architectures integrate compiler feedback through multi-agent systems employing Code Teachers reassessing bug addressing, Review Agents examining compilation logs generating structured prompts, reinforcement learning approaches with RLCF, PPOCoder, and COMPCODER, debugging agents using retrieval-augmented generation and Chain-of-Thought methodology, and tool-integrated approaches like CoderAgent [491, 393]. Domain-specific implementations include AutoChip incorporating compilation error messages, RTLFixer using retrieval-augmented generation for Verilog errors, and RTLCoder introducing code quality scoring [169].
Significant limitations persist in compiler feedback. Compilers report transformation failures through coarse error messages providing insufficient insight [493, 494, 117]. Project-specific context often cannot fit into LLM prompts, while workflow complexity creates inconsistent approaches [117, 493].
Accordingly, compiler feedback has evolved from basic validation to sophisticated iterative refinement providing detailed feedback including compilation success status, optimization pass information, and instruction counts [495, 38]. Effectiveness depends heavily on context—incorporating compiler feedback improves accuracy, though highlighting errors without resolution context shows limited improvement [117].
@@ 7.1.2 Static Analysis Feedback
LLMs demonstrate exceptional capabilities for understanding complex compiler messages and static analysis warnings [496, 497]. Correspondingly, static analysis tools analyze code without execution, employing predefined rules to enforce standards and detect bugs, though producing large volumes of warnings including false positives [498, 499, 500, 501, 502]. LLMs achieve 69-81% accuracy in prioritizing warnings and 81.13% precision with 94.64% recall in inspecting thousands of static warnings [503, 504].
Current systems employ comprehensive verification through multiple layers: syntactic correctness checking through compilers like GCC and Python interpreters, specialized tools including Black for formatting and nuXmv for formal verification, code complexity analysis utilizing OClint and srcSlice [505], and vulnerability detection employing Frama-C and Slither for Ethereum smart contracts [506, 507]. Static analysis extends to specification verification using Frama-C to verify AI-generated specifications [505].
Multi-agent architectures integrate static analysis into LLM-powered coding systems [124]. AutoSafeCoder implements three-agent frameworks with Coding Agents, Static Analyzer Agents, and Fuzzing Agents [138]. Reinforcement learning frameworks like REAL use program analysis-guided feedback combining automated signals with unit tests [508]. Applications span general programming detecting vulnerabilities in C [509], hardware description language generation with AutoChip [510], and HDLAgent supporting multiple languages [511].
@@ 7.1.3 Runtime Compilation Feedback
LLMs produce functional code but often struggle with correctness and optimization [512, 513, 253]. LLM-based coding agents leverage dynamic feedback from compilation and runtime processes to iteratively improve code output, transforming traditional one-shot code generation paradigms into iterative refinement processes [488, 514, 515, 516]. Coding agents utilize several feedback categories: binary pass/fail indicators, compilation feedback identifying syntax errors, failed tests feedback providing expected values with tracebacks, execution feedback providing runtime information, LDB feedback printing intermediate variable values, self-feedback determining compilation success, verification feedback checking unit tests and formal verification, and verbal feedback providing human-readable correction suggestions [243].
Several frameworks leverage compilation feedback: CompCoder implementing three-stage pipelines improving compilation success from 44.18% to 89.18%, CodeRL using actor-critic reinforcement learning with unit test signals [245, 279, 246], RLCF training pre-trained LLMs using feedback from grounding functions for code quality assessment [162], INTERVENOR using compiler feedback for agent collaboration, CoCoGen using static analysis to identify mismatches and iteratively fixing errors using compiler feedback and repository information [117], self-debugging frameworks like Cycle, LDB, and ReflectionCoder leveraging runtime execution information for iterative refinement [111, 163], CompilerGPT automating interaction between compilers, LLMs, and evaluation systems [515], and hardware-specific frameworks like AutoChip for RTL/Verilog generation demonstrating domain-specific effectiveness [511, 169].
Integration of compilation and runtime feedback finds applications across code translation ensuring semantic correctness through methods like TransMap and AlphaTrans [164], iterative code refinement where execution feedback serves as supervisory signal [165, 166, 167, 26], automatic program repair showing iterative patch creation benefits [168], visualization [167], and performance optimization where compilers provide structured feedback explaining optimization failures.
@@
SWE-Gym [299] 2024 2,438 verified instances (11 repos); 64,689 raw instances Constructs interactive execution environments from GitHub issues in 11 Python repos; each environment includes the full codebase and failing tests Repository-level RL environment Agents modify files and run unit tests within Docker; rewards are given when all tests pass, supporting iterative RL training
SWE-smith [310] 2025 50k tasks (250+ environments); 5,016 trajectories Automatically converts any GitHub repository into a SWE-gym environment, corrupting unit tests to generate interactive tasks Repository-level RL environment Environment provides file-localization and repair tools; agents interact through code edits and receive rewards when modified code passes tests
R2E-Gym [517] 2025 8.1k tasks (13 repos) Procedurally generates interactive RL environments from commit history using the SWE-Gen pipeline Commit/Repository-level environment Agents interact with full repositories via Docker; hybrid execution and static verifiers produce the reward signals
DeepSWE [518] 2025 $\approx$4.5k problems Builds a code RL environment by filtering R2E-Gym tasks to avoid overlap with evaluation benchmarks Repository-level RL environment Environment defines four actions (execute, search, edit, finish); reward is given only when all tests pass within a time limit
SWE-rebench [519] 2025 21,336 tasks Creates interactive environments from mined PRs with 1–15 file modifications and clear descriptions Repository-level interactive environment Uses a unified scaffold: agents edit code and run tests; environment rewards successful test passing, with 30% of repositories building successfully
Multi-turn SWE RL Env. [520] 2025 7,249 tasks Filtered subset of SWE-rebench tasks preprocessed into a multi-turn RL environment Repository-level multi-turn environment Agents can execute shell commands, edit files, search and submit patches; reward equals test success minus step penalties, encouraging efficient interactions
@@
COFFEE-GYM [521] 2024 Multiple problems (5 difficulty levels) Builds a feedback environment from human code submission histories with paired wrong/correct solutions Function-level feedback environment Agents generate feedback; the environment applies edits and runs $\sim$35.5 tests per problem to produce a reward
EvoCodeBench [522] 2024 275 samples (initial version) Curates repository-level generation tasks across 25 repos; environment updated every 6 months to avoid leakage Repository-level generation environment Agents must generate complete repositories; evaluation uses Pass@k and domain-specific metrics
DA-Code [523] 2024 $\approx$500 data analysis tasks Aggregates real-world data analysis tasks covering cleaning, ML, and exploration in interactive notebooks Notebook-level interactive environment Provides an executable sandbox for SQL/Python; evaluation based on correctness of analysis results
DS-1000 [524] 2024 1,000 problems Collects function-level tasks from seven data science libraries, each with prompt, solution, and tests Function-level offline tasks Agents write code to pass test_execution and test_string; limited interaction beyond code submission and test running
BioCoder [525] 2024 1,026 Python functions + 1,243 Java methods + 253 Rosalind problems Extracts bioinformatics code from GitHub and Rosalind projects, emphasizing cross-file dependencies Function-level environment with long contexts Evaluated using fuzz testing; environment requires handling long prompts (up to 2,600+ tokens) and cross-file reasoning
MLE-Dojo [526] 2025 200+ Kaggle tasks Encapsulates Kaggle competitions into an RL environment with unified data, evaluators, and leaderboards Competition-level interactive environment Agents can request information, validate and execute code; reward is based on relative performance via HumanRank
AgentPack [527] 2025 1.3M code edits Collects code edits co-authored by agents (Claude, Codex, Cursor) and humans from Git histories Commit-level offline dataset Offline supervised dataset with no interactive environment or defined RL reward

### v2

@@ 6.2.1 AI-Native Development Interfaces
The integration of coding agents into the developer’s workspace has transformed the traditional notion of an Integrated Development Environment (IDE). Instead of passively editing code, developers now engage in a continuous dialogue with the system. AI-native development interfaces embed large language models directly within the editor, combining code completion, conversational reasoning, and contextual memory [vaswani2017attention, devlin2019bert, brown2020language, raffel2019exploring].
Two dominant paradigms have emerged. The first emphasizes inline suggestion, where the system predicts the next tokens or statements directly within the editor, offering context-sensitive autocompletion and refactoring hints. The second paradigm emphasizes conversational interaction, in which the developer communicates with the agent using natural language to specify intent, request explanations, or perform debugging [sergeyuk2025human, sergeyuk2024ide, vaithilingam2023towards, robe2022designing, ross2023programmer, weber2024significant].
Studies show that integrating both paradigms within a unified interface produces higher task efficiency and user satisfaction. Inline completion supports rapid iteration, while conversational interaction enhances explainability and collaboration [sergeyuk2024ide, ross2023programmer]. Emerging IDEs such as Cursor and Q Developer integrate contextual memory and multi-turn reasoning, maintaining persistent dialogue histories to provide more coherent assistance [donato2025multimind, ross2023programmer, nam2023using, pinto2023developer].
@@ 6.2.2 Remote Development
The growth of cloud-native development practices has led to widespread adoption of remote and containerized development environments. Systems such as GitHub Codespaces [github_codespaces_2024] and DevContainer provide pre-configured, reproducible workspaces accessible via browsers or local editors. This paradigm allows coding agents to operate directly within standardized, resource-isolated environments synchronized with version control systems.
Remote development also supports computation-intensive workflows by allowing resource-demanding tasks to run on remote servers while the developer interacts through lightweight clients. SSH-based connections and browser-based terminals enable seamless transitions between local and cloud execution contexts. These capabilities are essential for agent-assisted programming, as they allow distributed collaboration while maintaining strict control over dependencies and runtime configurations [jetbrains_gateway_2022].
@@ 6.2.3 Tool Integration Protocol Standards
Standardization efforts have become increasingly important for achieving interoperability between coding agents and existing developer tools. The Model Context Protocol (MCP) [hou2025model, ray2025survey] defines a universal interface for exchanging contextual information such as source code, documentation, and environment state. The Language Server Protocol (LSP) [bork2023language, gunasinghe2021language] provides language-agnostic features including diagnostics, completion, and code navigation, while the Debug Adapter Protocol (DAP) [ernst2021deductive, garcia2024cross] standardizes debugging interactions.
Integrating these protocols allows coding agents to function as first-class participants within conventional development ecosystems. Agents can request, modify, and analyze project context seamlessly, while maintaining compatibility with existing IDEs and version control systems [li2025glue, koc2025mind, wu2025git]. This protocol convergence promotes an open, extensible environment for hybrid human–AI programming collaboration [gitkraken2025mcp, vs2025agent].
@@ 6.3.1 CI/CD Pipeline Integration
Continuous Integration and Continuous Deployment (CI/CD) frameworks are essential for maintaining the stability and quality of software systems. Within the context of coding agents, CI/CD processes ensure that automatically generated code passes through rigorous validation and testing stages before integration [singh2023microservices, dima2018waterfall, highsmith2002agile, beck2000extreme]. Pipeline-as-Code practices enable reproducibility by representing build and deployment workflows as versioned artifacts [singh2023microservices, ivanov2022extended]. In cloud-native ecosystems, Kubernetes and Jenkins-based automation frameworks such as KubeSphere facilitate modular CI/CD pipelines. These pipelines automatically handle code building, testing, and deployment, while ensuring that agent-generated contributions are subject to the same review and validation processes as human-written code.
In the emerging paradigm of Vibe Coding, CI/CD pipelines are further extended with agent participation. Agents generate code, perform static analysis, and conduct self-testing, but human oversight remains indispensable for critical verification stages. Designing CI/CD architectures that integrate agents safely and transparently is a key challenge for reliable large-scale deployment [singh2023microservices].
@@ 6.3.2 Cloud Compute Orchestration
Distributed orchestration platforms employ frameworks dynamically provisioning computational resources with production deployments demonstrating scale operating on 25,000 CPU core clusters organized into specialized pod configurations [feilhauer2016def]. The orchestration landscape has been enhanced by TOSCA (Topology and Orchestration Specification for Cloud Applications) providing interoperable models for describing cloud applications as typed directed topology graphs representing components as nodes and dependencies as links enabling developers to model cloud service structures with automated lifecycle management [bisicchia2023continuous, brogi2018tosker, dehury2021toscadata]. TOSCA’s platform-agnostic nature allows users to switch between cloud providers enabling applications to be migrated between vendors with minimal modification demonstrating strong compatibility with CI/CD technologies facilitating easier testing, re-deployment, and re-engineering [dehury2021toscadata, tomarchio2021torch], with modern platforms like FogArm providing autonomous adaptation capabilities responding to application specification changes from CI/CD pipelines and infrastructural variations detected through distributed monitoring [bisicchia2023continuous]. LLM-powered component integration adds complexity and opportunity where frameworks built on large language models, retrieval-augmented generation, and deep learning enable seamless integration with CI/CD pipelines supporting multi-format output generation and incorporating continuous learning for optimization [joshi2025architecting]. Modern implementations configure CI/CD pipelines to build, test, and deploy applications within containerized environments enabling faster build times and more frequent releases with Continuous Deployment integrating directly with LLMOps to automatically deploy model updates based on performance monitoring enabling automated model transitioning, deployment, and monitoring while maintaining high reliability using GitOps pipelines ensuring automatic redeployment of updated agent code or model weights with Canary releases and A/B testing [choi2025intelligent, song2025llm].
@@ 6.3.3 Multi-Agent Collaboration Frameworks
Frameworks such as AutoGen, CrewAI, MetaGPT, and LangGraph exemplify this paradigm. AutoGen organizes heterogeneous agents that perform role-specific tasks such as requirement analysis, implementation, and unit testing [wu2023autogen, zhu2023autogen, porsdam2023autogen]. CrewAI conceptualizes agent collaboration as a simulation of human teams, incorporating negotiation, task assignment, and feedback integration [taulli2025crewai]. MetaGPT introduces a meta-planning layer, where high-level agents design task hierarchies and delegate execution to subordinate agents, thereby achieving structured decomposition [hong2024metagpt, zhou2024metagpt]. LangGraph formalizes collaboration as a graph of nodes and edges representing agents and communication pathways, which enables systematic reasoning over dependencies and task flows [chen2025implementing].
These frameworks collectively redefine how coding agents can scale across complex development pipelines. They provide fault tolerance by redistributing tasks when individual agents fail, improve modularity through role specialization, and support iterative verification through feedback loops among agents. Recent studies suggest that multi-agent collaboration enhances both the reliability and interpretability of large-scale code synthesis [nasir2025code, williams2025multi].
Despite these advances, challenges remain regarding consistency, communication efficiency, and conflict resolution. Overly frequent message exchanges may lead to coordination overhead, while insufficient synchronization can cause semantic drift among agents. Ongoing research is exploring graph-structured communication protocols and shared reasoning buffers to improve coherence across collaborative agents [zeng2025edge, luo2025ai].
The development of multi-agent frameworks represents an essential direction for the evolution of Vibe Coding systems. These frameworks not only enhance scalability and adaptability but also embody a shift toward distributed intelligence in software engineering, where human developers and autonomous agents collaborate as peers within integrated ecosystems [hughes2025ai].
@@ 7.1.1 Syntax and Type Error Feedback
LLM-based coding agents rely on compiler feedback to iteratively improve generated code through error detection and repair cycles [han2025convcodeworld, peng2025criticlean, brown2020language, madaan2023self, bi2024iterative, bouzenia2024repairagent]. Compiler feedback encompasses multiple granularities: coarse-grained feedback provides binary compilation success, while fine-grained feedback delivers detailed error reasons and locations [ravi2025llmloop]. Beyond basic compilation feedback, systems integrate test feedback from automated frameworks and static analysis from tools like pylint [dai2025feedbackeval].
Modern architectures integrate compiler feedback through multi-agent systems employing Code Teachers reassessing bug addressing, Review Agents examining compilation logs generating structured prompts, reinforcement learning approaches with RLCF, PPOCoder, and COMPCODER, debugging agents using retrieval-augmented generation and Chain-of-Thought methodology, and tool-integrated approaches like CoderAgent [ravi2025llmloop, zhan2025coderagent]. Domain-specific implementations include AutoChip incorporating compilation error messages, RTLFixer using retrieval-augmented generation for Verilog errors, and RTLCoder introducing code quality scoring [tsai2023rtlfixer].
Significant limitations persist in compiler feedback. Compilers report transformation failures through coarse error messages providing insufficient insight [li2024review, siso2019evaluating, bi2024iterative]. Project-specific context often cannot fit into LLM prompts, while workflow complexity creates inconsistent approaches [bi2024iterative, li2024review].
Accordingly, compiler feedback has evolved from basic validation to sophisticated iterative refinement providing detailed feedback including compilation success status, optimization pass information, and instruction counts [grubisic2024compiler, huang2024opencoder]. Effectiveness depends heavily on context—incorporating compiler feedback improves accuracy, though highlighting errors without resolution context shows limited improvement [bi2024iterative].
@@ 7.1.2 Static Analysis Feedback
LLMs demonstrate exceptional capabilities for understanding complex compiler messages and static analysis warnings [abtahi2025augmenting, fang2024large]. Correspondingly, static analysis tools analyze code without execution, employing predefined rules to enforce standards and detect bugs, though producing large volumes of warnings including false positives [lacombe2023combining, srinivasarao2024software, li2024iris, keltek2025lsast, wadhwa2024core]. LLMs achieve 69-81% accuracy in prioritizing warnings and 81.13% precision with 94.64% recall in inspecting thousands of static warnings [becker2016effective, mohajer2023skipanalyzer].
Current systems employ comprehensive verification through multiple layers: syntactic correctness checking through compilers like GCC and Python interpreters, specialized tools including Black for formatting and nuXmv for formal verification, code complexity analysis utilizing OClint and srcSlice [tu2023isolating], and vulnerability detection employing Frama-C and Slither for Ethereum smart contracts [feist2019slither, hu2025qlpro]. Static analysis extends to specification verification using Frama-C to verify AI-generated specifications [tu2023isolating].
Multi-agent architectures integrate static analysis into LLM-powered coding systems [he2025llm]. AutoSafeCoder implements three-agent frameworks with Coding Agents, Static Analyzer Agents, and Fuzzing Agents [nunez2024autosafecoder]. Reinforcement learning frameworks like REAL use program analysis-guided feedback combining automated signals with unit tests [yao2025training]. Applications span general programming detecting vulnerabilities in C [dolcetti2024helping], hardware description language generation with AutoChip [chang2025data], and HDLAgent supporting multiple languages [kashanaki2025beyond].
@@ 7.1.3 Runtime Compilation Feedback
LLMs produce functional code but often struggle with correctness and optimization [tao2024codelutra, pearce2023examining, mathews2024test]. LLM-based coding agents leverage dynamic feedback from compilation and runtime processes to iteratively improve code output, transforming traditional one-shot code generation paradigms into iterative refinement processes [han2025convcodeworld, zheng2025vectrans, pirkelbauer2025compilergpt, li2024redo]. Coding agents utilize several feedback categories: binary pass/fail indicators, compilation feedback identifying syntax errors, failed tests feedback providing expected values with tracebacks, execution feedback providing runtime information, LDB feedback printing intermediate variable values, self-feedback determining compilation success, verification feedback checking unit tests and formal verification, and verbal feedback providing human-readable correction suggestions [dong2025survey].
Several frameworks leverage compilation feedback: CompCoder implementing three-stage pipelines improving compilation success from 44.18% to 89.18%, CodeRL using actor-critic reinforcement learning with unit test signals [le2022coderl, jiang2024survey, shojaee2023execution], RLCF training pre-trained LLMs using feedback from grounding functions for code quality assessment [jain2023coarse], INTERVENOR using compiler feedback for agent collaboration, CoCoGen using static analysis to identify mismatches and iteratively fixing errors using compiler feedback and repository information [bi2024iterative], self-debugging frameworks like Cycle, LDB, and ReflectionCoder leveraging runtime execution information for iterative refinement [ashrafi2025enhancing, ding2024cycle], CompilerGPT automating interaction between compilers, LLMs, and evaluation systems [pirkelbauer2025compilergpt], and hardware-specific frameworks like AutoChip for RTL/Verilog generation demonstrating domain-specific effectiveness [kashanaki2025beyond, tsai2023rtlfixer].
Integration of compilation and runtime feedback finds applications across code translation ensuring semantic correctness through methods like TransMap and AlphaTrans [ibrahimzada2024alphatrans], iterative code refinement where execution feedback serves as supervisory signal [gehring2024rlef, zeng2025acecoder, ni2025viscoder, yang2023intercode], automatic program repair showing iterative patch creation benefits [ruiz2025art], visualization [ni2025viscoder], and performance optimization where compilers provide structured feedback explaining optimization failures.
@@
SWE-Gym [pan2024training] 2024 2,438 verified instances (11 repos); 64,689 raw instances Constructs interactive execution environments from GitHub issues in 11 Python repos; each environment includes the full codebase and failing tests Repository-level RL environment Agents modify files and run unit tests within Docker; rewards are given when all tests pass, supporting iterative RL training
SWE-smith [yang2025swesmith] 2025 50k tasks (250+ environments); 5,016 trajectories Automatically converts any GitHub repository into a SWE-gym environment, corrupting unit tests to generate interactive tasks Repository-level RL environment Environment provides file-localization and repair tools; agents interact through code edits and receive rewards when modified code passes tests
R2E-Gym [jain2025r2e] 2025 8.1k tasks (13 repos) Procedurally generates interactive RL environments from commit history using the SWE-Gen pipeline Commit/Repository-level environment Agents interact with full repositories via Docker; hybrid execution and static verifiers produce the reward signals
DeepSWE [deepswe2025] 2025 $\approx$4.5k problems Builds a code RL environment by filtering R2E-Gym tasks to avoid overlap with evaluation benchmarks Repository-level RL environment Environment defines four actions (execute, search, edit, finish); reward is given only when all tests pass within a time limit
SWE-rebench [badertdinov2025swe] 2025 21,336 tasks Creates interactive environments from mined PRs with 1–15 file modifications and clear descriptions Repository-level interactive environment Uses a unified scaffold: agents edit code and run tests; environment rewards successful test passing, with 30% of repositories building successfully
Multi-turn SWE RL Env. [golubev2025training] 2025 7,249 tasks Filtered subset of SWE-rebench tasks preprocessed into a multi-turn RL environment Repository-level multi-turn environment Agents can execute shell commands, edit files, search and submit patches; reward equals test success minus step penalties, encouraging efficient interactions
@@
COFFEE-GYM [chae2024coffee] 2024 Multiple problems (5 difficulty levels) Builds a feedback environment from human code submission histories with paired wrong/correct solutions Function-level feedback environment Agents generate feedback; the environment applies edits and runs $\sim$35.5 tests per problem to produce a reward
EvoCodeBench [li2024evocodebench] 2024 275 samples (initial version) Curates repository-level generation tasks across 25 repos; environment updated every 6 months to avoid leakage Repository-level generation environment Agents must generate complete repositories; evaluation uses Pass@k and domain-specific metrics
DA-Code [huang2024code] 2024 $\approx$500 data analysis tasks Aggregates real-world data analysis tasks covering cleaning, ML, and exploration in interactive notebooks Notebook-level interactive environment Provides an executable sandbox for SQL/Python; evaluation based on correctness of analysis results
DS-1000 [lai2022natural] 2024 1,000 problems Collects function-level tasks from seven data science libraries, each with prompt, solution, and tests Function-level offline tasks Agents write code to pass test_execution and test_string; limited interaction beyond code submission and test running
BioCoder [tang2024biocoder] 2024 1,026 Python functions + 1,243 Java methods + 253 Rosalind problems Extracts bioinformatics code from GitHub and Rosalind projects, emphasizing cross-file dependencies Function-level environment with long contexts Evaluated using fuzz testing; environment requires handling long prompts (up to 2,600+ tokens) and cross-file reasoning
MLE-Dojo [qiang2025mle] 2025 200+ Kaggle tasks Encapsulates Kaggle competitions into an RL environment with unified data, evaluators, and leaderboards Competition-level interactive environment Agents can request information, validate and execute code; reward is based on relative performance via HumanRank
AgentPack [zi2025agentpack] 2025 1.3M code edits Collects code edits co-authored by agents (Claude, Codex, Cursor) and humans from Git histories Commit-level offline dataset Offline supervised dataset with no interactive environment or defined RL reward